Ma base bibliographique complète autour de l’intereferometrie radio et des methodes de deconvolution associees

A. Coulais

4 avril 2003

Références

Abl74 [1] J. Ables, Maximum Entropy spectral analysis, A&A Supplement Series, 15:383–393, 1974, . AMDT95 [2] N. Ageoges, J.-L. Monin, L. Desbat, and E. Tessier, Phase and image reconstruction from interferometric

imaging : integration of the phasors, A&A, Janvier 1995, .

Ali83 [3] C. Alissandrakis, Two-dimensional maps of the Sun with the Nançay Radioheliograph at 169 MHz, Rapport

interne, Décembre 1983, .

Arsac55 [4] J. Arsac, Nouveau réseau pour l’observation radioastronomique de la brillance du soleil à 9530 Mc/s, Comptes

Rendus de l’Académie des Sciences, 240:942–9445, 1955, .

AsBa94a [5] M. Aschwanden and T. Bastian, VLA stereoscopy of solar active regions. I. Method and tests, The Astrophysical

Journal, 426:425–433, Mai 1994, .

AsBa94b [6] M. Aschwanden and T. Bastian, VLA stereoscopy of solar active regions. II. Altitude, relative motion, and

center-to-limb darkening of 20 centimeter emission, The Astrophysical Journal, 426:434–448, Mai 1994, .

BaAl79 [7] E. Bajaja and G. van Albada, Complementing aperture synthesis radio data by short spacing compoments from

single dish observations, A&A, 75(0):251–254, 1979, .

Bart95 [8] D. Barthès, Cleaning power spectra of short, unevenly spaced time-series. application to LPVs, A&A Supplement

Series, 111(0):373–386, Juin 1995, .

Bat78 [9] R. Bates, Fringe visibility intensities may uniquely define brightness distributions, A&A, 70(0):L27–L29, mois

inconnu 1978, .

Bev93 [10] R. Bevensee, Maximum Entropy solutions to scientific problems, PTR Prentice Hall, Englewood Cliffs, New Jersey

07632, 1993.

Bhan78 [11] R. Bhandari, Maximum Entropy spectral analysis- some comments, A&A, 70(0):331–333, mois inconnu 1978, . Bir69 [12] Y. Biraud, A new approach for increasing the resolving power by data processing, A&A, 1(0):124–127, mois inconnu

1969, .

BiCh90 [13] Y. Biraud and A. Chevillot, Exact Fourier transform and interpolation of undersampled shifted autocorrelation

functions, A&A, 232(0):570–581, mois inconnu 1990, .

Blum72 [14] E. Blum, Les Radiotélescopes, no. 1454 in Que sais-je?, P.U.F., 1972. Boischot65 [15] A. Boischot, La Radioastronomie, no. 1182 in Que sais-je?, P.U.F., 1965.

BonK81 [16] J. Bonmartin and A. Kerdraon, Le système numérique de traitement de données du Radiohéliographe de

Nançay, L’onde Electrique, 0(0):11–16, Août 1981, .

Bra56a [17] R. Bracewell, Strip integration in radio astronomy, Australian Journal of Physics, 9:198–217, mois inconnu 1956,

.

Bra56b [18] R. Bracewell, Two-dimensional aerial smoothing in radio astronomy, Australian Journal of Physics, 9:297–314,

mois inconnu 1956, .

Bra65 [19] R. Bracewell, The Fourier Transform and it’s appplications, MacGraw-Hill,, mois inconnu 1965. HER79_Bra [20] R. Bracewell, Image reconstruction in Radio Astronomy, in Image reconstruction from projections, G. Herman,

ed., vol. 32 of Topics in applied physics, Springer Verlag, 1979, pp. 81–104.

Rob84_Bra [21] R. Bracewell, Inversion of nonplanar visibilities, in Indirect Imaging, J. Roberts, ed., URSI, Australie 1983,

1984, pp. 177–183.

Bra67 [22] R. Bracewell and A. Riddle, Inversion of fan-beam scans in radio astronomy, The Astrophysical Journal,

150(0):427–434, Novembre 1967, .

BraTh73 [23] R. Bracewell and A. Thompson, The main beam and ringlobes of an East-West rotation-synthesis array, The Astrophysical Journal, 182(0):77–94, Mai 1973, .

BrWal85 [24] R. Braun and R. Walterbos, A solution to the short spacing problem in radio interferometry, A&A, 143:307–312, mois inconnu 1985, .

Brem95 [25] M. Bremer, A 3D iterative deprojection technique. I. Development of the algorithm and tests, A&A Supplement Series, 112(0):551–566, mois inconnu 1995, .

BrSh84 [26] E. Brinks and W. Shane, A high resolution hydrogen line survey of Messier 31. I. Observations and data reduction, A&A Supplement Series, 55:179–186, Février 1984, .

Brouw71 [27] W. Brouw, Data processing for the Westerbork synthesis radio telescope, PhD thesis, Université de Leiden, Septembre 1971.

Buck91 [28] B. Buck and V. Macaulay, eds., Maximum Entropy in action, Oxford Science Publications, 1991.

HER79_BGH [29] T. Budinger, G. Gullberg, and R. Huesman, Emission computed tomography, in Image reconstruction from projections, G. Herman, ed., vol. 32 of Topics in applied physics, Springer Verlag, 1979, pp. 147–246.

Burg72 [30] J. Burg, The relationship between maximum entropy spectra and maximum likelihood spectra, Geophysics, 37(2):375–376, Avril 1972, .

Cap69 [31] J. Capon, High-resolution frequency-wavenumber spectrum analysis, Proc. IEEE, 57(8):1408–1418, Août 1969, .

Cap67 [32] J. Capon, R. Greenfield, and R. Kolker, Multidimensional Maximum–Likelihood processing of a large aperture seismic array, Proc. IEEE, 55(2):192–211, Février 1967, .

COB92 [33] M. Carbonell, R. Oliver, and J. Ballester, Power spectra of gapped time series: a comparison of several methods, A&A, 264(0):350–360, mois inconnu 1992, .

ChHo85 [34] W. Christiansen and J. Högbom, Radio Telescopes (second edition), Cambridge University Press, 1985.

Clar80 [35] B. Clark, An efficient implementation of the algorithm ”CLEAN”, A&A, 89(0):377–378, mois inconnu 1980, .

ThDad82_Clark [36] B. Clark, Large field mapping, in synthesis mapping, A. Thompson and L. D’Addario, eds., Proceedings of the NRAO-VLA workshop, Socorro, New Mexico, June 1982, National Radio Astronomy Observatory, Octobre 1983, pp. 10.1–10.18.

CPL85 [37] J. Clark, M. Palmer, and P. Lawrence, A transformation method for the reconstruction of functions from nonuniform spaced samples, IEEE Trans. on Acoustics, Speech, and Signal Processing, 33(4):1151–1165, Octobre 1985, .

Cole73 [38] T. Cole, Reconstruction of the image of a confined source, A&A, 24:41–45, mois inconnu 1973, .

Cogr89 [39] J. Combes, A. Grossmann, and P. Tchamitchian, eds., Wavelets, time-frequency methods and phase space, Inverse Problems and Theorical Imaging, Proceedings of the international conference, Marseille, France, December 1987, Sringer-Verlag, 1989.

ThDad82_Corn [40] T. Cornwell, Image restoration and the Clean technique, in synthesis mapping, A. Thompson and L. D’Addario, eds., Proceedings of the NRAO-VLA workshop, Socorro, New Mexico, June 1982, National Radio Astronomy Observatory, Octobre 1983, pp. 9.1–9.14.

Corn83 [41] T. Cornwell, A method of stabilizing the CLEAN algorithm, A&A, 121(0):281–285, mois inconnu 1983, .

PSB86_Corn [42] T. Cornwell, Deconvolution, in Synthesis Imaging in Radio Astronomy, R. Perley, F. Schaw, and A. Bridle, eds., vol. 6 of Whorkshop No. 13, NRAO, Green Bank, USA, 1986, ch. 7, pp. 109–122.

ZDN95_Corn [43] T. Cornwell, Imaging concepts, in Very long baseline interferometry and the VLBA, J. Zensus, P. Diamond, and P. Napier, eds., vol. 82 of Conference Series, Astronomical Society of the Pacific, 390 Asthon Avenue, San Francisco CA 94112, USA, 1995, ch. 3, pp. 39–56.

CoBr89 [44] T. Cornwell and R. Braun, Lecture 8: Deconvolution, in Synthesis Imaging in Radio Astronomy, R. Perley, F. Schaw, and A. Bridle, eds., vol. 6 of Conference Series, Astronomical Society of the Pacific, 390 Asthon Avenue, San Francisco CA 94112, USA, 1989, pp. 175–179.

CHU93 [45] T. Cornwell, M. Holdaway, and J. Uson, Radio–interferometric imaging of very large objects: implications for array design, A&A, 271(0):697–713, mois inconnu 1993, .

CoPe92 [46] T. Cornwell and R. Perley, Radio-interferometric imaging of very large fields. The problem of non-coplanar arrays, A&A, 261(0):353–364, mois inconnu 1992, .

CUH92 [47] T. Cornwell, J. Uson, and N. Haddad, Radio-interferometric imaging of spectral lines. The problem of continuum subtraction, A&A, 258(0):583–590, mois inconnu 1992, .

CoWi81 [48] T. Cornwell and P. Wilkinson, A new method for making maps with unstable radio interferometers, Monthly Notices of the Royal Astronomy Society, 196:1067–1086, mois inconnu 1981, .

CPLK93 [49] A. Coulais, J. Pinçon, F. Lefeuvre, and A. Kerdraon, Improvement of the Sun corona radio cartography, in Proceedings of contribution papers, F-2 section, physics of plasmas ans ionized media, Faculty of mathematics and physics, Charles University, Prague, Septembre 1993, pp. 18–24.

Davi94 [50] J. Davila, Solar tomography, The Astrophysical Journal, 423(0):871–877, Mars 1994, .

DaTh92 [51] J. Davila and W. Thompson, A rotating tomographic imager for solar extreme–ultraviolet/soft X–ray emission, The Astrophysical Journal, 389(0):L91–L93, Avril 1992, .

DaGr79 [52] M. Davison and F. Grünbaum, Convolution algorithms for arbitrary projections angles, IEEE Trans. on Nuclear Sciences, 26(2):2670–2673, Avril 1979, .

DeBr95 [53] A. Delaney and Y. Bresler, Multiresolution tomographic reconstruction using wavelets, IEEE Trans. of image processing, 4(6):799–813, Juin 1995, .

DeMa73 [54] B. Démidovitch and I. Marov, Eléments de calcul numérique, Editions MIR, 1973.

DoTh72 [55] G. Downs and A. Thompson, The distribution of linear polarization in Casiopeia A at wavelengths of 9.8 and 11.1 cm, The Astronomical Journal, 77(2):120–133, Mars 1972, .

Rob84_Dulk [56] G. Dulk, D. McLean, R. Manchester, and al, Image restoration techniques. how good are they?, in Indirect Imaging, J. Roberts, ed., URSI, Australie 1983, 1984, pp. 355–361.

ThDad82_Ekers [57] R. Ekers, Common image defects, in synthesis mapping, A. Thompson and L. D’Addario, eds., Proceedings of the NRAO-VLA workshop, Socorro, New Mexico, June 1982, National Radio Astronomy Observatory, Octobre 1983, pp. 12.1–12.18.

EBMP76 [58] I. El-Behery and R. MacPhie, Radio source parameters estimation by Maximum–Likelihood processing of variable baseline correlation interferometer data, IEEE Trans. on antennas and propagation, 24(2):163–173, Mars 1976, .

EBMP77 [59] I. El-Behery and R. MacPhie, Maximum–Likelihood estimation of source parameters from time–sampled outputs of a linear array, Journal of the Acoustical Society of America, 62(1):125–134, Juillet 1977, .

EBMP78 [60] I. El-Behery and R. MacPhie, Maximum likelihood estimation of the number, directions and strengths of point radio sources from variable baseline interferometer data, IEEE Trans. on antennas and propagation, 26(2):234–238, Mars 1978, .

EBMP80 [61] I. El-Behery and R. MacPhie, The resolution limit of a variable baseline synthetic aperture antenna, IEEE Trans. on antennas and propagation, 28(2):234–238, Mars 1980, .

MAX87_Escu [62] B. Escudié, Application de la représentation Temps-Fréquence à l’imagerie acoustique spatio–fréquentielle, in Méthodes et techniques de traitement du signal et applications aux mesures physiques, J. Max and collaborateurs, eds., vol. 2, Masson, Paris, 1987, ch. XXVI, pp. 332–360.

Foma73 [63] E. Fomalont, Earth-rotation aperture synthesis, Proc. IEEE, Special issue on radio and radar astronomy, 61(9):1211–1218, Septembre 1973, .

ThDad82_Foma [64] E. Fomalont, Image display, processing and analysis, in synthesis mapping, A. Thompson and L. D’Addario, eds., Proceedings of the NRAO-VLA workshop, Socorro, New Mexico, June 1982, National Radio Astronomy Observatory, Octobre 1983, pp. 11.1–11.39.

Fost95 [65] G. Foster, The cleanest Fourier spectrum, The Astronomical Journal, 109(4):1889, Avril 1995, .

Fost96 [66] G. Foster, Wavelets for period analysis of unevenly sampled time series, The Astronomical Journal, 112(4):1709– 1729, Octobre 1996, .

Fried72 [67] B. Frieden, Restoring with maximum likelihood and maximum entropy, Journal of the Optical Society of America, 62:511–518, 1972, .

GoHe74 [68] R. Gordon and G. Herman, Three–dimensional reconstruction from projections: a review of algorithms, International Journal of Cytology, 38:111–151, 1974, .

GGV95 [69] F. Gueth, S. Guilloteau, and F. Viallefond, A clean-based method for mosaic deconvolution, in YERAC 95, Cambridge, 1995.

HAYKIN85_Hay [70] S. Haykin, Radar array processing for angle of arrival estimation, in Array signal processing, S. Haykin, ed., Prentice-Hall, 1985, livre 4, pp. 194–292.

Her79 [71] G. Herman, ed., Image reconstruction from projections, vol. 32 of Topics in applied physics, Springer Verlag, 1979.

Hog74 [72] J. Högbom, Aperture synthesis with a non-regular distribution of interferometer baselines, Astron. Astrophys. Suppl., 15(0):417–426, mois inconnu 1974, .

Rob84_Hog [73] J. Högbom, Clean as a patern recognition procedure, in Indirect Imaging, J. Roberts, ed., URSI, Australie 1983, 1984, pp. 247–254.

HGGB77 [74] R. Huesman, G. Gullberg, W. Greenberg, and T. Budinger, RECLBL library users manual, Lawrence Berkeley laboratory, University of California, 1977.

Jaynes57 [75] E. Jaynes, Information theory and staitical mechanics, Phys. Rev., 106:620, 1957, .

Jer77 [76] A. Jerri, The Shannon sampling theorem. Its various extensions and applications: a tutorial review, Proc. IEEE, 65(11):1565–1596, Novembre 1977, .

HAYKIN85_Jus [77] J. Justice, Array processing in exploration seismology, in Array signal processing, S. Haykin, ed., Prentice-Hall, 1985, livre 2, pp. 6–114.

Kak84 [78] A. Kak, Image reconstruction from projections, in Digital Image Processing Techniques, M. Ekstrom, ed., vol. 2 of Computational Techniques, Academic Press, 1984, livre 4, pp. 111–170.

Kak85 [79] A. Kak, Tomographic imaging with diffracting an nondiffracting sources, in Array signal processing, J. Justice,

N. Owsley, J. Yen, and A. KAK, eds., Prentice-Hall, 1985, livre 6, pp. 351–428.

KaMa81 [80] S. Kay and S. Marple, Spectrum analysis : a modern perspective, Proc. IEEE, 69(11):1380–1419, Novembre 1981,

.

KerDel96 [81] A. Kerdraon and J.-M. Delouis, The Nançay Radioheliograph, in Coronal physics from radio and space observa-

tions, G. Trottet, ed., vol. 483, CESRA meeting, 3-7 Juin 1996, Nouan le Fuzelier, France, Avril 1997, pp. 192–201.

KiBo90 [82] S. Kim and N. Bose, Reconstruction of 2-D bandlimited discrete signals from nonuniform samples, Proc. IEEE,

137, Pt. F(3):197–204, Juin 1990, .

KNN81 [83] M. Komesaroff, R. Narayan, and R. Nityananda, the Maximum Entropy Method of image restoration-

properties ans limitations, A&A, 93:269–281, 1981, .

Kraus66 [84] J. Kraus, Radio Astronomy, McGraw-Hill book compagny, 1966. Kuhn82 [85] J. Kuhn, Recovering spectral information from unevenly sampled data: two machine-efficient solutions, The As-

tronomical Journal, 87(1):196–202, Janvier 1982, .

Lac83 [86] J.-L. Lacoume, Théorie du Signal, no. 2086 in Que sais-je?, P.U.F., 1983. MAX87_Lac [87] J.-L. Lacoume, Nouvelles méthodes d’analyse spectrale, in Méthodes et techniques de traitement du signal et

applications aux mesures physiques, J. Max and collaborateurs, eds., vol. 2, Masson, Paris, 1987, ch. XXI, pp. 131– 164.

Lagu86 [88] M. Lagunas, M. Santamaria, A. Gasull, and A. Moreno, Maximum likelihood filters in spectral estimation

problems, Signal Processing, 10(0):19–34, mois inconnu 1986, .

LAM97 [89] A. Lannes, E. Anterrieu, and P. Marechal, Clean and WIPE, A & A Supplement series, 123:183–198, Mai

1997, .

Lea91 [90] D. Leahy, Deprojection of emission in axially symmetric transparent systems, A&A, 247(0):584–589, mois inconnu

1991, .

Leg93 [91] C. Léger, R. Lédée, J. Thiel, L. Pourcelot, and F. Patat, Restitution complète d’un contour non fermé du

ventricule gauche par le modèle SFDP, journal inconnu, 0(0):0–0, mois inconnu 1993, .

Leq00 [92] J. Lequeux, Radio interferometry, in Coherent detection at millimeter wavelengths and their applications, E. et al.,

ed., centre de physique, Les Houches, mois inconnu 1990.

Lim84 [93] J. Lim, Image enhancement, in Digital Image Processing Techniques, M. Ekstrom, ed., vol. 2 of Computational

Techniques, Academic Press, 1984, livre 1, pp. 1–51.

LimMal81 [94] J. Lim and N. Malik, A new algorithm for two-dimensional maximum entropy power spectrum estimation, IEEE

Trans. on acoustics, speech, and signal processing, 29(3):401–413, Juin 1981, .

LoNa83 [95] A. Louis and F. Natterer, Mathematical problems of computerized tomography, Proc. IEEE, Special issue on

Tomography, 71(3):379–389, Mars 1983, .

LuMa92 [96] P. Luchini and F. Manzo, A fast conformal mapping algorithm with no FFT, Journal of Computational Physics,

101(0):368–374, mois inconnu 1992, .

Lucy92 [97] L. Lucy, Statistical limits to superresolution, A&A, 261(0):706–710, mois inconnu 1992, .

RH88 [98] C. Mercier, K.-L. Klein, and G. Trottet, The Nançay multifrequency Radioheliograph: new observations,

Advanced Space Research, 8(11):193–197, mois inconnu 1988, .

Mil92 [99] D. Milne, A simple method to enhance radio polarization feature, A&A, 261(0):694–696, mois inconnu 1992, . Mly93 [100] J. Mlynář, First computer tomography experimental results on the CASTOR tokamak, in Proceedings of contri-

bution papers, F-2 section, physics of plasmas and ionized media, Faculty of mathematics and physics, Charles University, Prague, Septembre 1993, pp. 177–182.

Mof68 [101] A. Moffet, Minimum–redundancy linear arrays, IEEE Trans. on antennas and propagation, 16(2):172–175, Mars

1968, .

Rob84_NN [102] R. Narayan and R. Nityananda, Maximum Entropy : Flexilibity versus fundamentalism, in Indirect Imaging,

J. Roberts, ed., URSI, Australie 1983, 1984, pp. 281–290.

NN86 [103] R. Narayan and R. Nityananda, Maximum Entropy image restoration in astronomy, Annual Revue of A&A,

24:127–170, 1986, .

New77 [104] W. Newman, A new method of multidimensional power spectral analysis, A&A, 54(0):369–380, mois inconnu 1977,

.

New78 [105] W. Newman, Notes on multidimensional power spectral analysis, A&A, 70(0):409–410, mois inconnu 1978, . Nil82 [106] R. Niland, Maximum reliable information obtainable by tomography, Journal of the Optical Society of America,

72(12):1677–1682, Décembre 1982, .

Rob84_Nil [107] R. Niland, Maximum reliable information in limited angle tomography, in Indirect Imaging, J. Roberts, ed., URSI, Australie 1983, 1984, pp. 325–331.

NN82 [108] R. Nityananda and R. Narayan, Maximum Entropy image reconstruction. a practical non-information theoretic approach, J. of Astrophys. Astr., 3:419–450, 1982, .

ONN94 [109] E. Otobe, J. Nakajima, K. Nishibori, T. Saito, and al., Two-dimensional direct images with a spatial FFT interferometer, Publication of the Astronomical Society of Japan, 46:503–510, mois inconnu 1994, .

Papa75 [110] G. Papadopoulos, A statistical technique for processing radio interferometer data, IEEE Trans. on antennas and propagation, 23(1):45–53, Janvier 1975, .

Papou62 [111] A. Papoulis, The Fourier Integral and its applications, MacGraw-Hill, 1962.

Papou66 [112] A. Papoulis, Error analysis in sampling theory, Proc. IEEE, 54:947–955, Juillet 1966, .

Papou77 [113] A. Papoulis, Signal analysis, MacGraw-Hill, 1977.

Park77 [114] R. Parker, Understanding inverse theory, Annual Revue of Earth and Planets Sciences, 5(0):35–64, mois inconnu 1977, .

ZDN95_Pear [115] T. Pearson, Non-imaging data analysis, in Very long baseline interferometry and the VLBA, J. Zensus, P. Diamond, and P. Napier, eds., vol. 82 of Conference Series, Astronomical Society of the Pacific, 390 Asthon Avenue, San Francisco CA 94112, USA, 1995, ch. 14, pp. 267–286.

PSB86 [116] R. Perley, F. Schaw, and A. Bridle, eds., Synthesis Imaging in Radio Astronomy, vol. 6 of Whorkshop No. 13, NRAO, Green Bank, USA, 1986.

PSB89 [117] , eds., Synthesis Imaging in Radio Astronomy, vol. 6 of Conference Series, Astronomical Society of the Pacific, 390 Asthon Avenue, San Francisco CA 94112, USA, 1989.

PGM96 [118] F. Peyrin, L. Garnero, and I. Magnin, Introduction à l’imagerie tomographique 2D et 3D reposant sur une propagation en ligne droite. Cas de la tomographie par rayon X par émission et par ultrasons, Traitement du signal, 13(4):380–450, Septembre 1996, .

PiLe91 [119] J.-L. Pinçon and f. Lefeuvre, Local characterization of homogeneous turbulence in a space plasma from simultaneous measurements of field components at several points in space, Journal of Geophysical Research, 96(A2):1789– 1802, Février 1991, .

Pon73 [120] Ponsonby, An Entropy measure of partially polarized radiation and its application to estimating radio sky polarization from incomplete aperture synthesis data by the maximum Entropy method, Monthly Notices of the Royal Astronomy Society, 163:369–380, 1973, .

PrWi90 [121] J. Prince and A. Willsky, Constrained sinogram restoration for limited-angle tomography, Optical Engineering, 29(5):535–544, Mai 1990, .

PrWi92 [122] J. Prince and A. Willsky, Hierarchical reconstruction unsing geometry and sinogram restoration, soumis à IEEE Trans. on image processing, 0, Février 1992, .

RH83 [123] T. Radioheliograph Group, The mark III Nançay Radioheliograph, Solar Physics, 88(0):383–390, mois inconnu 1983, .

RH89 [124] T. Radioheliograph Group, The mark IV Nançay Radioheliograph, Solar Physics, 120(0):193–204, mois inconnu 1989, .

RH93 [125] T. Radioheliograph Group, The Nançay multifrequency Radioheliograph: new developments and data acquisition for the solar physics community, Advanced Space Research, 13(9):411–414, mois inconnu 1993, .

NRH96 [126] T. Radioheliograph Group, Le Radiohéliographe de Nançay. Nouvelles perfomances : l’imagerie à deux dimensions, Novembre 1996.

Robert84 [127] J. Roberts, ed., URSI, Australie 1983, 1984.

HER79_Row [128] R. Rowland, Computer implementation of image reconstruction formulas, in Image reconstruction from projections, G. Herman, ed., vol. 32 of Topics in applied physics, Springer Verlag, 1979, pp. 9–80.

SNLIH81 [129] T. Sato, S. Norton, M. Linzer, O. Ikeda, and M. Hirama, Tomographic image reconstruction from limited projections using iterative revisions in image and transform spaces, Applied Optics, 20(3):395–399, Février 1981, .

SaNo95 [130] R. Sault and J. Noordam, Eliminating solar interference in radio interferometric spectrometry, A&A Supplement Series, 109(3):593–595, Mars 1995, .

Rob84_Schwab [131] F. Schwab, Optimal gridding of visibility in radio interferometry, in Indirect Imaging, J. Roberts, ed., URSI, Australie 1983, 1984, pp. 333–346.

Schwab84 [132] F. Schwab, Relaxing the isoplanation assumption in self–calibration; application to low–frequency radio interferometry, The Astronomical Journal, 89(0):1076–1081, Juillet 1984, .

Schwarz78 [133] U. Schwarz, Mathematical-statistical description of the iterative beam removing technic (Method CLEAN), A&A, 65(0):345–356, mois inconnu 1978, .

Rob84_Schwarz [134] U. Schwarz, The reliability of clean maps and the corrugation effect, in Indirect Imaging, J. Roberts, ed., URSI, Australie 1983, 1984, pp. 255–260.

SchWak91 [135] U. Schwarz and B. Wakker, Adding short spacings to synthesis maps in the ’sky’-domain, in Radio Interferometry : Theory, Technique and Applications, T. Cornwell and R. Perley, eds., vol. 19 of ASP Conference Series, IAU Collection 131, 1991.

SeFr78 [136] A. Segalovitz and B. Roy Frieden, A ”CLEAN”-type deconvolution algorithm, A&A, 70(0):335–343, mois inconnu 1978, .

Sha49 [137] C. Shannon and W. Weaver, The mathematical theory of communication, University of Illinois Press, Urbana, Il, 1949.

Shaw71 [138] L. Shaw, Spectral estimates from nonuniform samples, IEEE Trans. on Audio and Electroacoustics, 19(1):24–31, Mars 1971, .

Sou88 [139] M. Soumekh, Band-limited interpolation from unevenly spaced sampled data, IEEE Trans. on Acoustics, Speech, and Signal Processing, 36(1):110–122, Janvier 1988, .

PSB86_SS [140] R. Sramek and F. Schwab, Imaging, in Synthesis Imaging in Radio Astronomy, R. Perley, F. Schaw, and A. Bridle, eds., vol. 6 of Whorkshop No. 13, NRAO, Green Bank, USA, 1986, ch. 5, pp. 67–86.

SBLP94 [141] J. Stark, A. Bijaoui, B. Lopez, and C. Perrier, Image reconstruction by the wavelet transform applied to aperture synthesis, A&A, 283:349–360, Mars 1994, .

SDI84 [142] D. Steer, P. Dewdney, and M. Ito, Enhancement to the deconvolution algorithm ”CLEAN”, A&A, 137(0):159– 165, mois inconnu 1984, .

Swan82 [143] P. Swan, Discrete Fourier transforms of nonuniformly spaced data, The Astronomical Journal, 87(11):1608–1615, Novembre 1982, .

SwBa77 [144] W. Swindell and H. Barrett, Computerize tomography: taking sectional X rays, Physics Today, 0(0):32–41, Décembre 1977, .

ThBra74 [145] A. Thompson and R. Bracewell, Interpolation and Fourier transformation of fringe visibilities, The Astronomical Journal, 79(1):11–24, Janvier 1974, .

ThCr92 [146] A. Thompson and I. Craig, Automatic strategies for astrophysical inverse problems, A&A, 262(0):359–368, mois inconnu 1992, .

ThDad82 [147] A. Thompson and L. D’Addario, eds., synthesis mapping, Proceedings of the NRAO-VLA workshop, Socorro, New Mexico, June 1982, National Radio Astronomy Observatory, Octobre 1983.

TMS86 [148] A. Thompson, J. Moran, and G. Swenson, Interferometry and synthesis in Radio Astronomy, Wiley-Interscience, 1986.

UnAl96 [149] M. Unser and A. Aldroubi, A review of wavelets in biomedical applications, Proc. IEEE, Special issue on Wavelets, 84(4):626–638, Avril 1996, .

WakSch88 [150] B. Wakker and U. Schwarz, The multi-resolution CLEAN and its application to the short-spacing problem in interferometry, A&A, 200(0):312–322, mois inconnu 1988, .

WakSch91 [151] B. Wakker and U. Schwarz, The multi-resolution CLEAN, in Radio Interferometry : Theory, Technique and Applications, T. Cornwell and R. Perley, eds., vol. 19 of ASP Conference Series, IAU Collection 131, 1991.

Wer77 [152] S. Wernecke, Two-dimensional maximum entropy reconstruction of radio brightness, Radio science, 12(5):831– 844, Septembre 1977, .

Wild70 [153] J. Wild, De-convolution of barely resolved radio sources mapped with aerial beams of elleptical cross section, Australian Journal of Physics, 23:113–115, mois inconnu 1970, .

Yen56 [154] J. Yen, On the nonuniform sampling of bandwidth-limited signals, IRE Trans. Circuit Theory, CT–3:251–257, Décembre 1956, .

Yen85 [155] J. Yen, Image reconstruction in synthesis radio telescope array, in Array signal processing, J. Justice, N. Owsley, J. Yen, and A. KAK, eds., Prentice-Hall, 1985, livre 5, pp. 293–350.

ZDN95 [156] J. Zensus, P. Diamond, and P. Napier, eds., Very long baseline interferometry and the VLBA, vol. 82 of Conference Series, Astronomical Society of the Pacific, 390 Asthon Avenue, San Francisco CA 94112, USA, 1995.